# Constraints on Pre-Big-Bang Cosmology from Advanced LIGO and Advanced Virgo’s First Three Observing Runs

### v1

@@ 1 Introduction
The detection of gravitational waves (GWs) by Advanced LIGO [1] and Advanced Virgo [2] has ushered in a new era of observational astronomy [3, 4, 5]. GW observations have not only opened up a novel window to study the Universe, providing valuable insights into the physics of compact objects such as black holes and neutron stars, but have also served as a powerful tool to test the validity of general relativity in the strong-field regime. In addition to individual, high-amplitude GW events originating from merging compact binaries, the superposition of numerous weaker, unresolved GW signals can form a stochastic gravitational-wave background (SGWB). The study of SGWBs can yield crucial information about the properties and distribution of their sources, encompassing both astrophysical and cosmological origins [6, 7].
@@
Astrophysical contributions to the SGWB arise from a variety of sources, including merging compact binaries and core-collapse supernovae [8, 9]. On the other hand, cosmological sources are associated with various physical processes in the early Universe, such as cosmic phase transitions [10, 11, 12], scalar-induced GWs [13, 14, 15, 16, 17, 18, 19], cosmic strings [20, 21, 22, 23], cosmic domain walls [24, 25], and primordial density perturbations during inflation [26, 27, 28]. These cosmological sources are isotropic and predicted to generate SGWBs with distinct spectral features, which could provide valuable insights into the physics of the early Universe. However, it is important to note that these sources are not entirely beyond the framework of general relativity, which may limit our understanding of SGWBs in the context of quantum gravity.
@@
The primordial Universe, characterized by its extremely high energy scale, serves as a natural laboratory for studying quantum gravity. Among the existing theories of quantum gravity, string theory has garnered significant attention due to its potential to provide a unified description of all fundamental forces [29, 30]. String theory postulates that the fundamental building blocks of the Universe are tiny, vibrating strings of energy, which can give rise to the observed particles and forces [31, 32]. String cosmology, which applies the principles of string theory to the study of the early Universe, offers possible solutions to long-standing cosmological problems, such as the trans-Planckian problem [33] and the Big Bang singularity [34, 35]. One of the most iconic scenarios in string cosmology is the Pre-Big Bang (PBB) scenario [36, 37, 38, 39, 40, 41, 42, 43], which is based on the underlying duality symmetries of string theory and has emerged as a compelling framework for generating SGWBs [44]. This scenario predicts a cosmological phase of growing spacetime curvature and accelerated evolution, known as “superinflation” [45] followed by a non-singular transition to the standard radiation-dominated regime. As a result, a SGWB with a blue-tilted spectrum is naturally produced [46, 47]. This distinctive spectral shape could potentially distinguish the PBB scenario from other cosmological models, such as standard slow-roll inflation [48].
@@
Recent studies [49, 50] have investigated the compatibility of the PBB scenario with the stochastic signal detected by the North American Nanohertz Observatory for Gravitational Waves (NANOGrav) and concluded that the current formulation of the PBB model cannot adequately account for the observed data. These contrasting results highlight the need for further investigation into the viability of the PBB scenario in light of the latest observational evidence.
@@
In this paper, we will use data from the first three observing runs of the LIGO-Virgo collaboration to constrain the parameters of the PBB model. Although the LIGO-Virgo collaboration has not yet detected an SGWB signal, they have determined an upper limit to its amplitude, which enables us to constrain various cosmological models. The rest of this paper is organized as follows. In Section 2, we provide an overview of the PBB scenario and its predictions for the SGWB. In Section 3, we outline the methodology for obtaining model parameter constraints using data from Advanced LIGO and Advanced Virgo. Finally, we summarize our findings and discuss the implications of our results in Section 4.
2 SGWB from pre-big-bang cosmology
@@
In this section, we will briefly review the PBB scenario and its resulting SGWB. Due to the scale-factor duality of string cosmology [44], the evolution of our Universe should have a nearly mirror-symmetric phase of accelerated expansion preceding the decelerated expansion. This phase, referred to as the PBB scenario [36], provides a possible example of primordial tensor perturbations that peak at high frequencies and exhibit a blue-tilted spectrum at low frequencies. Here we review the derivation of SGWB. The spectral energy density of the SGWB present today inside our cosmic horizon can be written as:
$\Omega_{\textsc{gw}}(k,\tau_{0})=\frac{1}{\rho_{\rm crit}(\tau_{0})}\frac{d\rho_{\textsc{gw}}}{d\ln k},$ (2.1)
@@
$d\rho(\tau_{0})=2k\langle n_{k}(\tau_{0})\rangle\frac{d^{3}k}{8\pi^{3}}=\frac{k^{4}}{\pi^{2}}\langle n_{k}(\tau_{0})\rangle\ln k,$ (2.2)
where $\langle n_{k}(\tau_{0})\rangle$ is the number density of gravitons produced at $\tau_{0}$. We can obtain $\langle n_{k}(\tau_{0})\rangle$ by solving the following evolution equation for the tensor mode $h_{k}$ [40]:
$v_{k}^{\prime\prime}+\left(k^{2}-\frac{\xi^{\prime\prime}}{\xi}\right)v_{k}=0,$ (2.3)
where $v_{k}=\xi h_{k}$ and $\xi(\tau)$ is called the “pump field”. It can be seen from the above equation that it determines the dynamics of the fluctuation $h_{k}$. For the model chosen in this paper, the background is approximated as a sequence of five cosmic phases. The pumping field $\xi$ is a simple power-law behavior like $\xi=(M_{\text{Pl}}/\sqrt{2})|\tau/\tau_{1}|^{\alpha}$ in each phase, where $\tau_{1}$ denotes the time at the end of the string phase. Then the solution $h_{k}$ of Eq. (2.3) can be expressed by the first and second kinds of Hankel functions $H^{(1)}_{\nu}$, $H^{(2)}_{\nu}$ as
$h_{k}(\tau)=\left(\frac{2\tau_{1}}{M_{\text{Pl}}^{2}}\right)^{\frac{1}{2}}\left|\frac{\tau}{\tau_{1}}\right|^{\nu}\left[A_{+}(k)H_{\nu}^{(2)}(k\tau)+A_{-}(k)H_{\nu}^{(1)}(k\tau)\right].$ (2.4)
@@
$\langle n_{k}(\tau_{0})\rangle=\frac{4}{\pi}|A_{-}(k)|_{\tau=\tau_{0}}.$ (2.5)
By combining the above equation with Eqs. (2.1) and (2.2), we can get the the SGWB produced by the PBB scenario.
@@
The model considered in this paper is divided into five phases by four transition (at $\tau_{i}:\tau_{1},\tau_{\sigma},\tau_{d},\tau_{s}$). Here, $\tau_{\sigma},\tau_{d},\tau_{s}$ correspond to the beginning of a dust phase dominated by axion oscillations, the beginning of the post Big-Bang evolution, and the moment of transition from a low energy initial stage to a possible late attractor, respectively. In each of the above phases, the pump field has a simple power-law behavior. Specifically, in the initial PBB region $-\infty<\tau\leq-\tau_{1}$, we have two phases with the following pump field [40]
$\xi\sim\frac{M_{\text{Pl}}}{\sqrt{2}}(-\tau)^{1/2},~{}~{}~{}~{}~{}~{}~{}\tau<-\tau_{s};~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\xi\sim\frac{M_{\text{Pl}}}{\sqrt{2}}(-\tau)^{\beta-1},~{}~{}~{}~{}~{}~{}~{}-\tau_{s}<\tau<-\tau_{1},$ (2.6)
and in the subsequent bouncing state $-\tau_{1}<\tau<\tau_{0}$, the evolution of the universe is decelerated, and the pump field behavior of the corresponding three phases is:
$\xi\sim\frac{M_{\text{Pl}}}{\sqrt{2}}\tau,~{}~{}~{}-\tau_{1}<\tau<\tau_{\sigma};~{}~{}~{}~{}\xi\sim\frac{M_{\text{Pl}}}{\sqrt{2}}\tau^{2},~{}~{}~{}\tau_{\sigma}<\tau<\tau_{d};~{}~{}~{}~{}\xi\sim\frac{M_{\text{Pl}}}{\sqrt{2}}\tau,~{}~{}~{}\tau_{d}<\tau<\tau_{\text{0}}.$ (2.7)
Here $\beta$ is a parameter that describes the high-energy growth of the dilaton and the dynamics of internal dimensions. According to the specific forms of the pump field above, we can now represent the energy density fraction spectrum of SGWB as [40, 43]
@@
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\Omega_{\text{gw}}(f)=\left\{\begin{aligned} &\Omega_{\text{r}0}\left(\frac{H_{1}}{M_{\text{Pl}}}\right)^{2}\left(\frac{f_{d}}{f_{\sigma}}\right)^{2}\exp{[-(f-f_{1})/f_{1}]},~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}f>f_{1}\\ \\ &\Omega_{\text{r}0}\left(\frac{H_{1}}{M_{\text{Pl}}}\right)^{2}\left(\frac{f_{d}}{f_{\sigma}}\right)^{2}\left(\dfrac{f}{f_{1}}\right)^{3-|3-2\beta|},~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\qquad f_{\sigma}\lesssim f\lesssim f_{1}\\ \\ &\Omega_{\text{gw}}(f_{1})\left(\dfrac{f_{\sigma}}{f_{1}}\right)^{3-|3-2\beta|}\left(\dfrac{f}{f_{\sigma}}\right)^{1-|3-2\beta|},\qquad\quad\quad~{}~{}f_{d}\lesssim f\lesssim f_{\sigma}\\ \\ &\Omega_{\text{gw}}(f_{\sigma})\left(\dfrac{f_{d}}{f_{\sigma}}\right)^{1-|3-2\beta|}\left(\dfrac{f}{f_{d}}\right)^{3-|3-2\beta|},~{}~{}~{}~{}~{}~{}\qquad\,\,~{}f_{s}\lesssim f\lesssim f_{d}\\ \\ &\Omega_{\text{gw}}(f_{d})\left(\dfrac{f_{s}}{f_{d}}\right)^{3-|3-2\beta|}\left(\dfrac{f}{f_{s}}\right)^{3},\qquad\qquad\,\,\,\,\,~{}~{}~{}~{}~{}~{}~{}~{}f\lesssim f_{s}\end{aligned}\right.$ (2.8)
@@
where $f_{i}=1/(2\pi\tau_{i})$, $\Omega_{\text{r}0}\approx 4.15\times 10^{-5}h^{-2}$ is the current fraction of the energy density of radiation. For convenience, one can define three parameters as [43]
$z_{s}=\frac{\tau_{s}}{\tau_{1}}=\frac{f_{1}}{f_{s}},~{}~{}~{}~{}~{}~{}z_{\sigma}=\frac{\tau_{\sigma}}{\tau_{1}}=\frac{f_{1}}{f_{\sigma}},~{}~{}~{}~{}~{}~{}z_{d}=\frac{\tau_{d}}{\tau_{1}}=\frac{f_{1}}{f_{d}}.$ (2.9)
@@
$f_{1}=\frac{3.9\times 10^{11}}{2\pi}\left(\frac{H_{1}}{M_{\text{Pl}}}\right)^{1/2}\left(\frac{z_{\sigma}}{z_{d}}\right)^{1/2}\mathrm{Hz},$ (2.10)
@@
$\log_{10}\left(\frac{H_{1}}{M_{\text{Pl}}}\right)$ $=$ $\frac{2}{5-n_{\rm s}}\left\{\log_{10}\left[\frac{4.2\pi^{2}}{T^{2}(H_{1})}\right]-9+(1-n_{\rm s})(\log_{10}1.5-27)\right.$ (2.11)
$+\left.(1-n_{\rm s}-2\beta)\log_{10}z_{s}+\frac{n_{\rm s}-1}{2}\left(\log_{10}z_{\sigma}-\log_{10}z_{d}\right)\right\}.$
Here, $n_{\text{s}}=0.9649\pm 0.0042$ [51], and
$T(H_{1})\approx 0.13\left(\frac{H_{1}}{M_{\text{Pl}}}\right)^{1/6}z_{d}^{1/4}z_{\sigma}^{-\frac{7}{12}}+0.25\left(\frac{H_{1}}{M_{\text{Pl}}}\right)^{-1/6}z_{d}^{-1/4}z_{\sigma}^{\frac{7}{12}}-0.01.$ (2.12)
@@
Now, the spectrum of SGWB (2.8) is determined by only four undetermined parameters: $\beta,z_{s},z_{\sigma}$, and $z_{d}$. It is worth pointing out that the spectrum (2.8) can be fitted by the following functions [43]
$\Omega_{\text{GW}}(f)$ $=$ $\Omega_{\text{PBB}}f^{3}(f^{2}+f_{s}^{2})^{-\frac{|3-2\beta|}{2}}(f^{2}+f_{d}^{2})^{-1}(f^{2}+f_{\sigma}^{2})(f^{2}+f_{1}^{2})^{\frac{|3-2\beta|-3}{2}}$ (2.13)
×$\times\exp\left(\arctan\frac{f}{f_{1}}-\frac{f}{f_{1}}\right).$
In next section, we will use data from LIGO-Virgo observations to constrain these parameters by searching for the PBB signal.
3 Data analysis

### v2

# Pre-Big-Bang Cosmology Meets Gravitational Wave Astronomy: Constraints from Advanced LIGO and Advanced Virgo’s First Three Observing Runs

@@ I Introduction
The detection of gravitational waves (GWs) by Advanced LIGO Aasi et al. (2015) and Advanced Virgo Acernese et al. (2015) has ushered in a new era of observational astronomy Abbott et al. (2019, 2021a, 2023). GW observations have not only opened up a novel window to study the Universe, providing valuable insights into the physics of compact objects such as black holes and neutron stars, but have also served as a powerful tool to test the validity of general relativity in the strong-field regime. In addition to individual, high-amplitude GW events originating from merging compact binaries, the superposition of numerous weaker, unresolved GW signals can form a stochastic gravitational-wave background (SGWB). The study of SGWBs can yield crucial information about the properties and distribution of their sources, encompassing both astrophysical and cosmological origins Regimbau (2011); Christensen (2019).
@@
Astrophysical contributions to the SGWB arise from a variety of sources, including merging compact binaries and core-collapse supernovae Zhu et al. (2011); Crocker et al. (2017). On the other hand, cosmological sources are associated with various physical processes in the early Universe, such as cosmic phase transitions Kibble (1980); Witten (1984); Mazumdar and White (2019), scalar-induced GWs Ananda et al. (2007); Baumann et al. (2007); Garcia-Bellido et al. (2016); Inomata et al. (2017); Garcia-Bellido et al. (2017); Kohri and Terada (2018); Cai et al. (2019); Papanikolaou et al. (2024), cosmic strings Kibble (1976); Sarangi and Tye (2002); Damour and Vilenkin (2005); Siemens et al. (2007), cosmic domain walls Vilenkin and Everett (1982); Sikivie (1982), and primordial density perturbations during inflation Starobinsky (1979); Turner (1997); Bar-Kana (1994). These cosmological sources are isotropic and predicted to generate SGWBs with distinct spectral features, which could provide valuable insights into the physics of the early Universe. However, it is important to note that these sources are not entirely beyond the framework of general relativity, which may limit our understanding of SGWBs in the context of quantum gravity.
@@
The primordial Universe, characterized by its extremely high energy scale, serves as a natural laboratory for studying quantum gravity. Among the existing theories of quantum gravity, string theory has garnered significant attention due to its potential to provide a unified description of all fundamental forces Witten (1995); Aharony et al. (2000). String theory postulates that the fundamental building blocks of the Universe are tiny, vibrating strings of energy, which can give rise to the observed particles and forces Schwarz (1982); Kaplunovsky (1985). String cosmology, which applies the principles of string theory to the study of the early Universe, offers possible solutions to long-standing cosmological problems, such as the trans-Planckian problem Martin and Brandenberger (2001) and the Big Bang singularity Borde and Vilenkin (1994); Borde et al. (2003). One of the most iconic scenarios in string cosmology is the pre-big bang (PBB) scenario Gasperini and Veneziano (1993); Gasperini et al. (1997); Gasperini and Veneziano (2016); Fan and Zhu (2008); Gasperini (2016); Li et al. (2019); Gasperini (2022); Tan et al. (2025a); Ben-Dayan et al. (2024), which is based on the underlying duality symmetries of string theory and has emerged as a compelling framework for generating SGWBs Veneziano (1991). This scenario predicts a cosmological phase of growing spacetime curvature and accelerated evolution, known as “superinflation” Lucchin and Matarrese (1985) followed by a non-singular transition to the standard radiation-dominated regime. As a result, a SGWB with a blue-tilted spectrum is naturally produced Brustein et al. (1995); Jiang et al. (2023). This distinctive spectral shape could potentially distinguish the PBB scenario from other cosmological models, such as standard slow-roll inflation Linde (1991).
@@
Recent studies Tan et al. (2025b); Conzinu et al. (2025) have investigated the compatibility of the PBB scenario with the stochastic signal detected by the North American Nanohertz Observatory for Gravitational Waves (NANOGrav) and concluded that the current formulation of the PBB model cannot adequately account for the observed data. These contrasting results highlight the need for further investigation into the viability of the PBB scenario in light of the latest observational evidence. While previous studies Tan et al. (2025b); Conzinu et al. (2025) focused on testing the PBB model against pulsar timing array (PTA) observations in the nanohertz frequency band, the present work extends this analysis to the significantly higher frequency band accessible to ground-based GW detectors. This multi-band approach is essential for comprehensively testing the PBB scenario, as the predicted SGWB spectrum spans many orders of magnitude in frequency, with different frequency ranges probing different epochs of the PBB evolution.
@@
In this paper, we will use data from the first three observing runs of the LIGO-Virgo collaboration to constrain the parameters of the PBB model. Although the LIGO-Virgo collaboration has not yet detected an SGWB signal, they have determined an upper limit to its amplitude, which enables us to constrain various cosmological models. The rest of this paper is organized as follows. In Section II, we provide an overview of the PBB scenario and its predictions for the SGWB. In Section III, we outline the methodology for obtaining model parameter constraints using data from Advanced LIGO and Advanced Virgo. Finally, we summarize our findings and discuss the implications of our results in Section IV.
II SGWB from pre-big-bang cosmology
@@
In this section, we will briefly review the PBB scenario and its resulting SGWB. Due to the scale-factor duality of string cosmology Veneziano (1991), the evolution of our Universe should have a nearly mirror-symmetric phase of accelerated expansion preceding the decelerated expansion. This phase, referred to as the PBB scenario Gasperini and Veneziano (1993), provides a possible example of primordial tensor perturbations that peak at high frequencies and exhibit a blue-tilted spectrum at low frequencies. Here we review the derivation of SGWB. The spectral energy density of the SGWB present today inside our cosmic horizon can be written as:
$\Omega_{\textsc{gw}}(k,\tau_{0})=\frac{1}{\rho_{\rm crit}(\tau_{0})}\frac{d\rho_{\textsc{gw}}}{d\ln k},$ (1)
@@
$d\rho(\tau_{0})=2k\langle n_{k}(\tau_{0})\rangle\frac{d^{3}k}{8\pi^{3}}=\frac{k^{4}}{\pi^{2}}\langle n_{k}(\tau_{0})\rangle\ln k,$ (2)
where $\langle n_{k}(\tau_{0})\rangle$ is the number density of gravitons produced at $\tau_{0}$. We can obtain $\langle n_{k}(\tau_{0})\rangle$ by solving the following evolution equation for the tensor mode $h_{k}$ Gasperini (2016):
$v_{k}^{\prime\prime}+\left(k^{2}-\frac{\xi^{\prime\prime}}{\xi}\right)v_{k}=0,$ (3)
where $v_{k}=\xi h_{k}$ and $\xi(\tau)$ is called the “pump field”. It can be seen from the above equation that it determines the dynamics of the fluctuation $h_{k}$. For the model chosen in this paper, the background is approximated as a sequence of five cosmic phases. The pumping field $\xi$ is a simple power-law behavior like $\xi=(M_{\text{Pl}}/\sqrt{2})|\tau/\tau_{1}|^{\alpha}$ in each phase, where $\tau_{1}$ denotes the time at the end of the string phase. Then the solution $h_{k}$ of Eq. (3) can be expressed by the first and second kinds of Hankel functions $H^{(1)}_{\nu}$, $H^{(2)}_{\nu}$ as
$h_{k}(\tau)=\left(\frac{2\tau_{1}}{M_{\text{Pl}}^{2}}\right)^{\frac{1}{2}}\left|\frac{\tau}{\tau_{1}}\right|^{\nu}\left[A_{+}(k)H_{\nu}^{(2)}(k\tau)+A_{-}(k)H_{\nu}^{(1)}(k\tau)\right].$ (4)
@@
$\langle n_{k}(\tau_{0})\rangle=\frac{4}{\pi}|A_{-}(k)|_{\tau=\tau_{0}}.$ (5)
By combining the above equation with Eqs. (1) and (2), we can get the the SGWB produced by the PBB scenario.
[…]
[width=0.35]figure.eps
Figure 1: The time evolution of the Hubble horizon $H^{-1}$ (dashed line) and the scale factor $a(\tau)$ (solid line) in the PBB model. The shaded areas represent the causal connection spatial profiles of the Hubble size $H^{-1}$ at various epochs. Different $\tau$ represent the time when different phases began.
@@
The model considered in this paper is divided into five phases by four transition (at $\tau_{i}:\tau_{1},\tau_{\sigma},\tau_{d},\tau_{s}$). Here, $\tau_{1},\tau_{\sigma},\tau_{d},\tau_{s}$ correspond to the time at the end of the string phase, the beginning of a dust phase dominated by axion oscillations, the beginning of the post Big-Bang evolution, and the moment of transition from a low energy initial stage to a possible late attractor, respectively. In Fig. 1, we present the evolution of the Hubble radius $H^{-1}$ and the scale factor $a(\tau)$ in the PBB model as they evolve with conformal time. We also mark the approximate times $\tau_{1},\tau_{\sigma},\tau_{d},\tau_{s}$ on the graph. In each of the above phases, the pump field has a simple power-law behavior. Specifically, The specific form of $\xi$ is Gasperini (2016)
@@
$\xi\sim\left\{\begin{aligned} \hskip-2.84544pt&\frac{M_{\text{Pl}}}{\sqrt{2}}(-\tau)^{1/2},~~~~~\tau<-\tau_{s}\\ &\frac{M_{\text{Pl}}}{\sqrt{2}}(-\tau)^{\beta-1},~~~-\tau_{s}<\tau<-\tau_{1}\\ &\frac{M_{\text{Pl}}}{\sqrt{2}}\tau,~~~~~~~~~~~~~-\tau_{1}<\tau<\tau_{\sigma}\\ &\frac{M_{\text{Pl}}}{\sqrt{2}}\tau^{2},~~~~~~~~~~~~\tau_{\sigma}<\tau<\tau_{d}\\ &\frac{M_{\text{Pl}}}{\sqrt{2}}\tau,~~~~~~~~~~~~~\tau_{d}<\tau<\tau_{\text{0}}.\end{aligned}\right$. (6)
The parameter $\beta$ has deep theoretical roots in string cosmology. From the definition $\beta\equiv\frac{\mathrm{d}\log g_{s}}{\mathrm{d}\log a}$, it represents the rate of growth of the four-dimensional string coupling $g_{s}$ with respect to the scale factor $a$. This coupling evolution is crucial for the transition from the dilaton-dominated phase to the string phase and eventually to the post-big bang era. According to the specific forms of the pump field above, we can now represent the energy density fraction spectrum of SGWB as Gasperini (2016); Ben-Dayan et al. (2024)
$\Omega_{\text{GW}}(f)=\left\{\begin{aligned} \hskip-2.84544pt&\Omega_{\rm PBB}\exp{[-(f-f_{1})/f_{1}]},~~~~f>f_{1}\\ &\Omega_{\rm PBB}\left(\dfrac{f}{f_{1}}\right)^{\beta_{1}},~~~~~~~~~~~~~~~~f_{\sigma}\lesssim f\lesssim f_{1}\\ &\Omega_{\text{gw}}(f_{1})\left(\dfrac{f_{\sigma}}{f_{1}}\right)^{\beta_{1}}\left(\dfrac{f}{f_{\sigma}}\right)^{\beta_{2}},~f_{d}\lesssim f\lesssim f_{\sigma}\\ &\Omega_{\text{gw}}(f_{\sigma})\left(\dfrac{f_{d}}{f_{\sigma}}\right)^{\beta_{2}}\left(\dfrac{f}{f_{d}}\right)^{\beta_{1}},~f_{s}\lesssim f\lesssim f_{d}\\ &\Omega_{\text{gw}}(f_{d})\left(\dfrac{f_{s}}{f_{d}}\right)^{\beta_{1}}\left(\dfrac{f}{f_{s}}\right)^{3}.~~~f\lesssim f_{s}\end{aligned}\right$. (7)
Here $f_{i}=1/(2\pi\tau_{i})$, $\beta_{1}=3-|3-2\beta|$, and $\beta_{2}=1-|3-2\beta|$. The dimensionless amplitude $\Omega_{\text{PBB}}$ is given by
$\Omega_{\text{PBB}}=\Omega_{\text{r}0}\left(\frac{H_{1}}{M_{\text{Pl}}}\right)^{2}\left(\frac{f_{d}}{f_{\sigma}}\right)^{2},$ (8)
@@
where $\Omega_{\text{r}0}\approx 4.15\times 10^{-5}h^{-2}$ is the critical fraction of the current radiant energy density. For convenience, one can define three parameters as Ben-Dayan et al. (2024)
$z_{s}=\frac{\tau_{s}}{\tau_{1}}=\frac{f_{1}}{f_{s}},~~~~~~z_{\sigma}=\frac{\tau_{\sigma}}{\tau_{1}}=\frac{f_{1}}{f_{\sigma}},~~~~~~z_{d}=\frac{\tau_{d}}{\tau_{1}}=\frac{f_{1}}{f_{d}}.$ (9)
@@
$f_{1}=\frac{3.9\times 10^{11}}{2\pi}\left(\frac{H_{1}}{M_{\text{Pl}}}\right)^{1/2}\left(\frac{z_{\sigma}}{z_{d}}\right)^{1/2}\mathrm{Hz},$ (10)
@@
$\log_{10}\left(\frac{H_{1}}{M_{\text{Pl}}}\right)$ $=$ $\frac{2}{5-n_{\rm s}}\left\{\log_{10}\left[\frac{4.2\pi^{2}}{T^{2}(H_{1})}\right]-9\right$. (11)
$+\left.(1-n_{\rm s})(\log_{10}1.5-27)\right$.
$+\left.(1-n_{\rm s}-2\beta)\log_{10}z_{s}\right$.
$+\left.\frac{n_{\rm s}-1}{2}\left(\log_{10}\frac{z_{\sigma}}{z_{d}}\right)\right\}.$
Here, $n_{\text{s}}=0.9649\pm 0.0042$ Aghanim et al. (2020), and
$T(H_{1})\approx$ $0.13\left(\frac{H_{1}}{M_{\text{Pl}}}\right)^{1/6}z_{d}^{1/4}z_{\sigma}^{-\frac{7}{12}}$ (12)
$+0.25\left(\frac{H_{1}}{M_{\text{Pl}}}\right)^{-1/6}z_{d}^{-1/4}z_{\sigma}^{\frac{7}{12}}-0.01.$
@@
Now, the spectrum of SGWB (7) is determined by only four undetermined parameters: $\beta,z_{s},z_{\sigma}$, and $z_{d}$. It is worth pointing out that the spectrum (7) can be fitted by the following functions Ben-Dayan et al. (2024)
$\Omega_{\text{GW}}(f)$ $=$ $\Omega_{\text{PBB}}f^{3}(f^{2}+f_{s}^{2})^{-\frac{|3-2\beta|}{2}}(f^{2}+f_{d}^{2})^{-1}$ (13)
$\times(f^{2}+f_{\sigma}^{2})(f^{2}+f_{1}^{2})^{\frac{|3-2\beta|-3}{2}}$
$\times\exp\left(\arctan\frac{f}{f_{1}}-\frac{f}{f_{1}}\right).$
It is important to note that different frequency ranges of the spectrum probe different physical aspects of the PBB scenario. The low-frequency branch ($f<f_{s}$) with its $f^{3}$ behavior is determined by the initial conditions and early dynamics. The intermediate branches ($f_{s}<f<f_{1}$) with their $\beta$-dependent power laws directly probe the dynamics during the high-curvature string phase. The high-frequency cutoff region ($f\sim f_{1}$) constrains the bounce transition scale. This multi-band structure makes it essential to test the PBB model across many orders of magnitude in frequency, from nHz (PTA) to Hz-kHz (LIGO-Virgo) bands. Figure 2 illustrates representative PBB spectra for different values of $\beta$, demonstrating how the predicted signal amplitudes compare to the current sensitivity of LIGO-Virgo detectors. In next section, we will use data from LIGO-Virgo observations to constrain these parameters by searching for the PBB signal.
[…]
[width=0.5]ogw.pdf
Figure 2: Energy density spectrum of the SGWB produced by PBB scenarios compared to LVK sensitivity. The blue, orange, and green curves show representative PBB spectra with $\beta=-0.19$, $\beta=-0.08$, and $\beta=0.08$, respectively. For all three cases, we fix the other parameters to $\log_{10}z_{s}=10$, $\log_{10}z_{d}=5$, and $\log_{10}z_{\sigma}=4.6$. The gray shaded region represents the power-law integrated sensitivity from the LIGO-Virgo first three observing runs.
III Data analysis